OpenStudy (anonymous):

Which means "the sum of w and 3.4 is greater than or equal to 10.5"?

OpenStudy (ja1):

Ok do you know what "the sum of" would mean?

OpenStudy (anonymous):

I want to say that it is adding something

OpenStudy (anonymous):

I want to say that it is this 3.4 + w ≥ 10.5

OpenStudy (ja1):

You thought pretty well ;)

OpenStudy (anonymous):

hahaha thanks

OpenStudy (ja1):

@cz031081 Oh and welcome to OpenStudy! If you need help with anything feel free to message me or tag me if you see an issue :)

OpenStudy (anonymous):

The distributive property of multiplication justifies which of the following statements? 4(y – z) = 4(y) + 4(-z) 3(s) + 3(t) = 9(st) 8 + (-8) = 0 (a + b) + c = (c + a) + b

OpenStudy (anonymous):

thanks I appreciate it

OpenStudy (ja1):

Well let's try to eliminate the ones that are simply bogus, can you pick out the ones that are not even distributing?

OpenStudy (anonymous):

so I am guessing that the last one justifies the distributive property

OpenStudy (ja1):

And no problem I'm here to help :)

OpenStudy (anonymous):

second and third one

OpenStudy (ja1):

Actually no, the last one justifies a property, just not distributive, it represents the transitive property, those parenthesis are tricky huh? Remember that to distribute basically you do this: (i used bogus variables as placeholders)